OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + y = 6 6x + 2y = 8 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

3x + y = 6 + 6x + 2y = 8 __________________________ First pick one variable to solve for x or y 9x = 14 x=14/9 Then you can just plug in x into both equations to get y If you get 2 different answers for y, then there is no solution

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

@**izzy** Since when is y + 2y = 0?